### v1

# Approximate predictive control barrier function for discrete-time systems

Abstract
We propose integrating an explicit approximation of a predictive control barrier function (PCBF) in a safety filter framework. The approximated PCBF is implicitly defined through an optimal control problem and allows guaranteeing invariance of an implicitly defined safe set as well as stability of this safe set within a larger domain of attraction. By extending existing theoretical analysis of the PCBF, we establish inherent robustness of the original algorithm and translate the guarantees to input-to-state stability of the proposed algorithm with respect to possible approximation errors, recovering the same guarantees in the absence of approximation errors. The proposed algorithm allows certifying inputs with respect to state constraint satisfaction through a single function evaluation and filtering unsafe inputs through a control barrier function based safety filter, which is independent of the time horizon of the original predictive optimisation problem, resulting in significant online computational benefits. We demonstrate the stability properties of the proposed algorithm on a linear system example as well as its use a fast safety filter for miniature race cars in simulation.
{IEEEkeywords}
Constrained control, machine learning, NL predictive control, stability of nonlinear systems
1 INTRODUCTION
Safety filters provide a modular mechanism which enables certifying, and if necessary modifying, any control input such that safety in terms of constraint satisfaction can be guaranteed. The filters are designed to be agnostic with respect to the main control performance goal and therefore aim to minimise the distance between the proposed input and the filtered input which is applied to the system. Thereby, safety filters are suitable tools to enhance the safety properties of various systems. The filters can be used to augment human agents operating systems, such as, e.g., automotive vehicles as demonstrated on miniature race cars in [1], and have also demonstrated their effectiveness in complex robotics applications, such as, e.g., walking on stepping stones with a legged robot in [2].
The basic principle of safety filters consists in obtaining a control invariant set within the state constraints, such that inputs applied to the system guarantee that the system state remains within this invariant set. In order to obtain such a control invariant set, three main branches have been pursued, as discussed extensively in [3]. Hamilton-Jacobi reachability analysis, see, e.g., [4], allows computing the maximal control invariant subset to constraints, however the method suffers from poor computational scalability with respect to increasing state dimension. Control barrier functions (CBFs), see, e.g., [5] for an overview, aim to alleviate the computational restrictions, by explicitly computing, potentially smaller, control invariant sets through either expert knowledge or more scalable computational methods compared to the Hamilton-Jacobi reachability, such as semi-definite programming or sum-of-squares optimisation, see, e.g., [6]. Finally, predictive safety filters have been proposed in [7]. They employ model predictive control (MPC) techniques, see, e.g., [8], in order to implicitly define a control invariant set through an online predictive optimisation problem. These predictive methods make use of discrete time dynamics formulations and allow to easily provide recursive feasibility guarantees as well as extensions due to the well developed MPC theory to, e.g., robust treatments of disturbances, stochastic disturbances in [9] or using online data for uncertain systems to improve safety filter performance in [10].
In practical systems, considering hard constraints in predictive control methods often leads to infeasibility due to unmodeled dynamics or sensor errors. Therefore, the state constraints on the system are often softened, as proposed, e.g., in [11] and used in [1], by introducing a slack variable. However, this leads in general to a loss of theoretical properties as soon as the system leaves the constraints. In [12], a constraint softening mechanism is proposed, which is able to provide theoretical guarantees when constraints are violated. Connections of this mechanism to CBFs are investigated and it is shown that by minimising a function which depends on the slack variables, a control barrier function is implicitly defined. This result implies that the implicitly defined safe set of a predictive safety filter is stable within a larger domain of attraction, such that the system is able to recover from unexpected disturbances or model mismatch. The main drawback of the method in [12] is the need to solve two predictive optimisation problems in order to guarantee stability, which results in a computationally demanding algorithm.
1.1 Contributions
In this work, we extend initial results in [13] and aim to reduce the computational overhead of the predictive optimisation based PCBF method in [12] by obtaining an explicit approximation of the implicitly defined CBF. The explicit approximation can be used in a safety filter formulation similarly as in control barrier function based safety filters, such that only a single time step needs to be predicted using a decrease constraint on the value of the control barrier function. This decrease constraint further enables the introduction of a hyperparameter, which allows tuning the rate of stability decrease in closed-loop based on desired system behaviour. In comparison, in [12] the system is required to achieve the maximal possible decrease towards the implicitly defined safe set at every time step. We further extend the theoretical results in [12], by showing a rigorous connection between the control barrier function definition in [12] and more widely used class $\mathcal{K}$ comparator function based definitions. These relations allow to show that the method in [12] benefits from inherent robustness properties, see, e.g., [14], which are crucial for deployment of predictive control methods in real world applications. By using the insights from the inherent robustness of the algorithm proposed in [12], input-to-state stability of the proposed approximation-based algorithm is established with respect to the approximation errors. This result provides a theoretical understanding of how approximation errors in the formulation potentially impact the closed-loop performance of the system. In addition, we propose a continuous extension of the PCBF, which is negative within the implicitly defined safe set. Thereby, approximation errors lead to a significant reduction in safety filter interventions when the system is safe compared to approximating the original PCBF, which is 00 within the safe set. Finally, we show how the explicit approximation can be used in order to certify inputs using a single function evaluation, resulting in large computational speedups when the main controller proposes safe inputs.
1.2 Related work
Safety filters have been extensively researched and a review of state-of-the-art methods can be found in [3]. The first main branch of research consists of control barrier function based safety filters, originally proposed in [15] for continuous time systems. The work in [16] proposed discrete-time control barrier functions. However, the proposed formulation uses a barrier function which approaches infinity as the state approaches the boundary of the safe set, which is not suitable for investigating the stability of the safe set within a larger domain. In [17], a backup CBF formulation is proposed, which allows using a backup controller to enhance an invariant safe set, by computing gradients of a CBF on the closed-loop trajectory of the backup controller. Thereby, any state which reaches the safe set under the backup controller is still considered safe, and the safe set is effectively augmented. The predictive safety filters in [18] and [12] show relations to this method, where the optimiser is free to choose all control inputs in the prediction horizon, rather than selecting them with a prior fixed policy. In [19], the predictive safety filter is augmented with a stability guarantees of the origin using a Lyapunov decrease constraint. Compared to these works, we combine the advantages of the existence of inputs which stabilise a safe set and render it invariant, while using a single CBF decrease constraint rather than a full prediction horizon, reducing the computational effort. Due to the discrete-time formulation of the predictive controllers, the resulting CBF optimisation problem is a NLP rather than a QP in [17], however the prediction horizon is not actually included in the online optimisation, effectively reducing the number of constraints. The proposed formulation in [20] uses a look-ahead to determine time instances, where given a path of the system, a CBF constraint is violated. Thereby, the control actions are modified such that safety can be guaranteed. However, this method does not consider the case when the input is constrained. Finally, [21] proposes a reachability-based safety filter formulation in continuous time, which makes use of a finite lookahead time and establishes connections to model predictive safety filter methods. However, the method uses numerical approximations of the gradients of the CBF through sampling to impose the CBF decrease and it is unclear how approximation errors can impact the closed-loop performance. In comparison, we provide a rigorous analysis of how approximation errors of the learned CBF impact the closed-loop behaviour of the system.
Another main branch of research for safety filters is Hamilton-Jacobi reachability analysis, see, e.g., [4]. In [22], connections between viscosity solution of the Hamilton-Jacobi-Isaacs variational inequality, retrieving the maximum robust control invariant subset to given constraints and CBFs are established and it is shown how this solution can be used in a quadratic program. This result is extended in [23] using a forward reachability approach, which shows that any CBF is also a forward reachability solution within the maximum robust control invariant set. Compared to these results, we consider stability of the control invariant set implicitly defined through the predictive safety filter optimisation problem within a larger domain in a soft-constrained manner. Furthermore, it holds that if the prediction horizon of the predictive safety filter goes to infinity, the maximum control invariant set is also recovered, see, e.g., [24, Definition 2.9].
Numerous works consider approximation methods to reduce the computational drawbacks of online predictive control methods. Works such as, e.g., [25, 26, 27] are able to guarantee closed-loop properties such as stability and robust constraint satisfaction with respect to input uncertainty. However, the safety filter policy is in general non-continuous and multi-valued, such that continuous approximations can lead to arbitrary approximation errors. Other works consider learning the optimal value function of an MPC, such as, e.g., [28, 29] to increase performance rather than to address general safety considerations. An overview for learning safety certificates is provided in [30] with a main focus on certificate synthesis for continuous-time systems. In [31], a control barrier function is learned from expert demonstrations with theoretical guarantees in case of Lipschitz continuity. These results could in principle be applicable for the proposed approximation scheme, however Lipschitz continuity of the predictive control barrier function is in general not guaranteed. In [32], a state-action control barrier function is approximated which makes use of contractivity of a safe set to ensure invariance of the resulting safe set for small enough approximation errors with a convex online optimisation problem. This results in the need to sample over the state-action space compared to sampling over the state space as proposed in this paper. Additionally, stability of the safe set itself is not considered in [32] and recursive feasibility only holds if the approximation error is small enough, whereas the proposed formulation ensures feasibility everywhere.
[…]
2 PRELIMINARIES
Consider the discrete-time, nonlinear dynamical system
$x(k+1)=f(x(k),u(k))\;\;\forall k\in\mathbb{N},\$ (1)
with state $x\in\mathbb{R}^{n}$ and input $u\in\mathbb{R}^{m}$ and continuous dynamics $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$. The system is subject to state and input constraints
$\begin{split}x(k)&\in\mathcal{X}:=\{x\in\mathbb{R}^{n}\mid c_{x}(x)\leq 0\},\\ u(k)&\in\mathcal{U},\end{split}$ (2)
where the input constraints are considered to arise from, e.g., physical actuator limitations, while the state constraints arise from system specifications, such as, e.g., a limit on the velocity of a car on a highway lane. In practical predictive optimisation based methods, such constraints are typically formulated as soft constraints, see, e.g., [11], in order to preserve feasibility under unknown disturbances of the optimisation problems which are solved online. Accordingly, we define the soft-constrained set
$\mathcal{X}(\xi):=\{x\in\mathbb{R}^{n}\mid c_{x}(x)\leq\xi\},$
where $\xi\in\mathbb{R}^{n_{x}}$ is a slack variable. The constraint sets satisfy the following assumption.
Assumption 1 (Constraint Sets)
The constraint sets $\mathcal{U}$ and $\mathcal{X}(\xi)$ are compact for all $0\leq\xi<\infty$. Furthermore, it holds that $c_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n_{x}}$ is continuous.
The goal is to design a predictive safety filter for system (1). Given an initial state that lies outside the constraints $\mathcal{X}$, the system states should converge back into $\mathcal{X}$ and an initial state within a safe set should remain therein. In order to ensure such closed-loop stability and constraint satisfaction, we make use of the concept of control barrier functions. These functions typically define an invariant set, see, e.g. [33], as a zero sublevel set and provide stability guarantees within some domain with respect to this invariant set. Therefore, they are often considered to be an extension of a classical Lyapunov function, where the origin is stabilised. CBFs have been extensively studied in continuous-time dynamical systems, see, e.g., [5] for an overview, but have also been proposed in discrete-time in [16]. In this work, we consider two definitions for discrete-time CBFs introduced in [12] and [8]. These definitions differ in the properties the CBF needs to satisfy and we establish their relationship in Section 4 in order to analyse robust stability of the PCBF and the proposed algorithm.
Definition 1 (Control Barrier Function - Continuous [12])
A function $h:\mathcal{D}\to\mathbb{R}$ is a discrete-time control barrier function with safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}\subset\mathcal{D}\subset\mathbb{R}^{n}$ if the following conditions hold:
  1. (1)
    $\mathcal{S},\mathcal{D}$ are compact and non-empty sets,
  2. (2)
    $h(x)$ is continuous on $\mathcal{D}$,
  3. (3)
    $\exists\Delta h:\mathcal{D}\to\mathbb{R}$, with $\Delta h$ continuous and $\Delta h(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$ such that:
$\hskip 17.07182pt\forall x\in\mathcal{D}\setminus\mathcal{S}:$
$\inf_{u\in\mathcal{U}}\{h(f(x,u))\mid f(x,u)\in\mathcal{D}\}-h(x)\leq-\Delta h(x),$ (3a)
$\hskip 17.07182pt\forall x\in\mathcal{S}\>:\inf_{u\in\mathcal{U}}h(f(x,u))\leq 0.$ (3b)
This first definition requires that the CBF $h$ is continuous and defines a set $\mathcal{S}$ as a zero sublevel set which is control invariant through the condition (3b). Additionally, for any state outside the safe set, i.e., any $x\in\mathcal{D}\setminus\mathcal{S}$, there exists a continuous, positive decrease $\Delta h(x)$ of the CBF value. Inputs which satisfy the CBF conditions are considered safe to be applied to the given dynamical system.
Definition 2 (Safe inputs)
Given a CBF $h$ according to Definition 1, an input $u$ applied to system (1) at state $x$ is said to be safe, if $u\in\mathcal{U}_{h}(x)\subseteq\mathcal{U}$ with
$\mathcal{U}_{h}(x)\coloneqq\left\{\begin{array}[]{ll}U_{1}(x),&\text{if }x\in\mathcal{D}\setminus\mathcal{S},\\ U_{2}(x),&\text{if }x\in\mathcal{S},\end{array}\right.$ (6)
where $U_{1}(x)\coloneqq\{u\in\mathcal{U}\mid h(f(x,u))-h(x)\leq-\Delta h(x)\}$ and $U_{2}(x)\coloneqq\{u\in\mathcal{U}\mid h(f(x,u))\leq 0\}$.
In order to formalise the closed-loop property of applying safe inputs to a system, we make use of the notion of asymptotic stability of invariant sets.
Definition 3 (Asymptotic stability [12])
Let $\mathcal{S}$ and $\mathcal{D}$ be non-empty, compact and positively invariant sets for the system $x(k+1)=f(x(k),\kappa(x(k))$ with $\kappa:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ such that $\mathcal{S}\subset\mathcal{D}$. The set $\mathcal{S}$ is called an asymptotically stable set in $\mathcal{D}$, if for all $x(0)\in\mathcal{D}$, the following conditions hold:
$\forall\epsilon>0\;\exists\delta>0:$ $\;\|x(0)\|_{\mathcal{S}}<\delta\Rightarrow\forall k>0:\|x(k)\|_{\mathcal{S}}<\epsilon,$
$\exists\bar{\delta}:$ $\;\|x(0)\|_{\mathcal{S}}<\bar{\delta}\Rightarrow\lim_{k\rightarrow\infty}x(k)\in\mathcal{S}$
Applying safe inputs to system (1) ensures that the set $\mathcal{S}$ is asymptotically stable in the domain $\mathcal{D}$ as given by the following result.
Theorem 1 ([12])
Let $\mathcal{D}\subset\mathbb{R}^{n}$ be a non-empty and compact set. Consider a control barrier function $h:\mathcal{D}\rightarrow\mathbb{R}$ according to Definition 1 with $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}\subset\mathcal{D}$. If $\mathcal{D}$ is a forward invariant set for system (1) under $u(k)=\kappa(x(k))$ for all $\kappa:\mathcal{D}\rightarrow\mathcal{U}$ with $\kappa(x)\in\mathcal{U}_{h}(x)$, then it holds that
  1. 1.
    $\mathcal{S}$ is a forward invariant set
  2. 2.
    $\mathcal{S}$ is asymptotically stable in $\mathcal{D}$
The second definition of a control barrier function which we consider is given in [8, Appendix B.39] and makes use of class $\mathcal{K}$ comparator functions [^1] A function $\alpha:[0,c]\rightarrow\mathcal{R}$ is class $\mathcal{K}$ if it is continuous, strictly increasing, $\alpha(0)=0$..
Definition 4 (Control barrier function - Class $\mathcal{K}$ [8])
A function $h:\mathcal{D}\rightarrow\mathbb{R}$ is a discrete-time control barrier function with safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)=0\}\subset\mathcal{D}$ if the following hold:
  1. (1)
    $\mathcal{S},\mathcal{D}$ are compact and non-empty sets,
  2. (2)
    $\exists\;\alpha_{1},\alpha_{2}\in\mathcal{K}$, such that $\alpha_{1}(\|x\|_{\mathcal{S}})\leq h(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}})$,
  3. (3)
    $\exists\;\alpha_{3}\in\mathcal{K}$ such that
$\inf_{u\in\mathcal{U}}\{h(f(x,u))|\;f(x,u)\in\mathcal{D}\}-h(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}})$ (8)
A definition based on such comparator functions is advantageous due to the ease of extending stability guarantees if there are additional exogeneous disturbances acting on the system in (1). In Section 4, we will show that a CBF fulfilling Definition 1 implies existence of a CBF according to Definition 4. A similar relation of a continuous Lyapunov decrease implying the existence of a class $\mathcal{K}$ decrease is shown in [34], but uses class $\mathcal{K}_{\infty}$ bounds on the original Lyapunov function rather than the continuity of $h$ in Definition 1. The established relation between the two CBF definitions allows showing inherent robustness properties of the algorithm in [12]. Furthermore, it enables the input-to-state stability guarantees of the approximate predictive CBF approach proposed in this paper.
Remark 1
In [8, Appendix B.39], a function $h$ which fulfils Definition 4 is denoted a control Lyapunov function, however in the scope of this paper we distinguish a control Lyapunov function as function which fulfils Definition 4 with $\mathcal{S}=\{0\}$ from a CBF with a general compact set $\mathcal{S}$.
3 PREDICTIVE CONTROL BARRIER FUNCTIONS
In this section, we review the concept of predictive safety filters, as proposed originally in [7] and adapted in [12], and how they can be related to the concept of a control barrier function (CBF). At every time step, a controller proposes a potentially unsafe control action, denoted as $u_{\mathrm{p}}(k)$. Such a controller could be optimising a primary control objective, such as a robot operating in a cluttered environment or even a human operating a car. The aim of the safety filter is to certify, whether the proposed input is safe to apply to the system in terms of the state and input constraints (2), and modify it if necessary. In order to guarantee constraint satisfaction for all times, the predictive safety filter makes use of model predictive control techniques. The predictive safety filter problem proposed in [12] which is solved online at every time step is given as
$\min_{u_{i|k}}\;\;$ $\lvert\lvert{u_{0|k}-u_{\mathrm{p}}(k)}\rvert\rvert$ (9a)
s.t. $\forall\,i=0,\dots,N-1,$
$x_{0|k}=x(k),$ (9b)
$x_{i+1|k}=f(x_{i|k},u_{i|k}),$ (9c)
$u_{i|k}\in\mathcal{U},$ (9d)
$x_{i|k}\in\overline{\mathcal{X}}_{i}(\xi^{*}_{i|k}),$ (9e)
$h_{f}(x_{N|k})\leq\xi^{*}_{N|k},$ (9f)
Given the current state $x(k)$, a trajectory of states $x_{i|k}$ and inputs $u_{i|k}$ is planned according to the system dynamics in (9c) such that the state and input constraints (2) are satisfied for all predicted time steps. We note that as the initial state can lie outside the state constraints $\mathcal{X}$, the state constraints are formulated in terms of soft constraints with respect to slack variables $\xi^{*}_{i|k}$ which are computed according to a second optimisation problem discussed later in this section. The state constraints are subject to an additional tightening, i.e., $\overline{\mathcal{X}}_{i}(\xi_{i|k}):=\{x\in\mathbb{R}^{n}\mid c_{x}(x)\leq-\Delta_{i}+\xi_{i|k}\}$ with $\Delta_{i}-\Delta_{i-1}>\Delta_{0}=0$, in order to establish stability guarantees of the closed-loop. A terminal constraint (9f) is imposed on the final predicted state, where $h_{f}$ is a CBF according to Definition 1. As the explicit formulation of $h_{f}$ is used in the optimisation problem, it is assumed to be known.
Assumption 2 (Terminal control barrier function)
A terminal control barrier function $h_{f}:\mathcal{D}_{f}\rightarrow\mathbb{R}$ with domain $\mathcal{D}_{f}:=\{x\in\mathbb{R}^{n}\mid h_{f}(x)\leq\gamma_{f}\}$ and corresponding safe set $\mathcal{S}_{f}:=\{x\in\mathbb{R}^{n}\mid h_{f}(x)\leq 0\}\subseteq\overline{\mathcal{X}}_{N-1}(0)$ exists according to Definition 1 for the system dynamics (1) and is known.
A number of design procedures exist to design such a CBF, including Lyapunov-based methods such as semi-definite programming for quadratic functions, e.g., in [35], which can also be applied to nonlinear systems as discussed, e.g., in [12, Section IV], and sum-of-squares optimisation for polynomial systems and Lyapunov functions in, e.g., [36]. While obtaining a control barrier function for nonlinear systems is a challenging problem in general, we note that a continuous, local Lyapunov function around an equilibrium point is sufficient to obtain $h_{f}$ satisfying Definition 1 as discussed in [12, Section IV] and conservativeness introduced by small terminal sets $\mathcal{S}_{f}$ can be counteracted by increasing the prediction horizon $N$ in the optimisation problem (9).
In order to be minimally invasive, given the proposed performance input $u_{\mathrm{p}}(k)$ at any given time step, the safety filter problem minimises the distance of the first predicted input $u_{0|k}$ and $u_{\mathrm{p}}(k)$. Trivially, if the proposed performance input $u_{\mathrm{p}}(k)$ allows constructing a safe backup trajectory, it is also the optimal solution to the safety filter problem. At every time step, after obtaining a solution to (9), the input $u(k)=u^{*}_{0|k}$ is applied to the system.
[…]
Building upon the introduced predictive safety filter formulation, it is shown in [12], that a similar enhancement mechanism can be formulated for CBFs to increase the domain of attraction of the closed-loop system. As in practical applications, the initial state $x(0)$ may not be feasible without soft constraints or unmodelled external disturbances could result in infeasible states for the optimisation problem, a recovery mechanism is required, which ensures that the system converges back to an invariant subset of the constraints $\mathcal{X}$ in such events. Through their property of stability of invariant sets in a given domain, CBFs offer a theoretical framework which is suitable for such a recovery mechanism. This enhancement property for the domain of attraction of the closed-loop system is achieved by computing the slack variables $\xi^{*}_{i|k}$ in (9) according to the following optimisation problem, which minimises the state constraint violations and a weighted penalty on the terminal slack $\xi_{N|k}$, i.e.
$h_{\mathrm{PB}}(x(k)):=\min_{u_{i|k},\xi_{i|k}}\;\;$ $\alpha_{f}\xi_{N|k}+\sum_{i=0}^{N-1}\lvert\lvert{\xi_{i|k}}\rvert\rvert$ (10a)
s.t. $\forall\,i=0,\dots,N-1,$
$x_{0|k}=x(k),$ (10b)
$x_{i+1|k}=f(x_{i|k},u_{i|k}),$ (10c)
$u_{i|k}\in\mathcal{U},$ (10d)
[…]
The optimal value function of (10), i.e. $h_{\mathrm{PB}}(x)$, is then shown to be a CBF according to Definition 1 in [12].
Theorem 2 ([12])
Let Assumptions 1 and 2 hold. The minimum (10) exists and if $\alpha_{f}$ is chosen sufficiently large, then the optimal value function $h_{\mathrm{PB}}(x(k))$ defined in (10a) is a (predictive) control barrier function according to Definition 1 with domain $\mathcal{D}_{\mathrm{PB}}:=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)\leq\alpha_{f}\gamma_{f}\}$ and safe set $\mathcal{S}_{\mathrm{PB}}:=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)=0\}$.
Invariance of the safe set $\mathcal{S}_{\mathrm{PB}}$ in Theorem 2 results from recursive feasibility of the problem with $\xi^{*}_{i|k}=0$ from standard arguments, while the existence of a continuous, positive decrease $\Delta h_{\mathrm{PB}}$ in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ is ensured through the terminal CBF $h_{f}$ in Assumption 2 if $\xi^{*}_{N|k}>0$ and through the additional tightening $\Delta_{i}$ of the state constraints $\overline{\mathcal{X}}_{i}(\xi_{i|k})$ if $\xi^{*}_{N|k}=0$. Clearly, if $\xi^{*}_{i|k}=0$, the problem (9) is solved with additional tightening on the state constraints and the corresponding tightenings $\Delta_{i}$ should therefore be chosen small.
4 MAIN THEORETICAL ANALYSIS
In this section, we provide the main theoretical analysis which provides the foundation of the approximation algorithm in Section 5.1. The contributions are as follows. We first show that any CBF according to Definition 1 implies existence of a CBF according to Definition  4. This result enables analysis of the closed-loop behaviour of the predictive CBF algorithm with respect to possible errors. These errors can be in a first step considered to be exogeneous disturbances, implying inherent robustness, as defined in [14], of the original algorithm. In a second step, these errors will be considered to be errors when approximating the optimal value function $h_{\mathrm{PB}}$, enabling a principled analysis of the impact of regression errors on the closed-loop behaviour of the system.
4.1 Relation of CBF Definitions
In order to show the existence of a CBF according to Definition 4 given a CBF according to Definition 1, we need to relate the continuous function $h_{\mathrm{PB}}$ and its corresponding decrease $\Delta h_{\mathrm{PB}}$ to class $\mathcal{K}$ comparator functions. For the sake of exposition, we therefore first provide the class $\mathcal{K}$ lower bound and its proof, initially proposed in [13], in a slightly altered manner.
Lemma 1 (reformulated from [13])
Consider the compact and non-empty sets $\mathcal{S}$ and $\mathcal{D}$, with $\mathcal{S}\subset\mathcal{D}$ and a continuous function $\gamma:\mathcal{D}\setminus\mathcal{S}\rightarrow\mathbb{R}_{\geq 0}$, with $\gamma(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$. Then there exists a $\mathcal{K}$-function $\alpha:[0,\bar{r}]\rightarrow\mathbb{R}_{\geq 0}$ with $\bar{r}:=\max_{x\in\mathcal{D}}\|x\|_{\mathcal{S}}$ such that
$\gamma(x)\geq\alpha(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\mathcal{S}$ (11)
Proof 4.3.
Define the auxiliary function $p\!:\![0,\overline{r}]{\to}\mathbb{R}_{\geq 0}$ as
$p(r)\!:=\!\begin{cases}0&\text{if }r=0,\\ &\\ \parbox{133.72786pt}{\footnotesize$\min\limits_{y}\;\;\gamma(y)\\ \text{s.t.}\;\;y\in\mathcal{D}\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}<r\}$}&\mathrm{otherwise}.\end{cases}$ (12)
Since $\gamma$ is a continuous function and $\mathcal{D}\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}<r\}$ is a compact set, the above minimum is attained for all $r\in(0,\overline{r}]$. It holds that
$p(\|x\|_{\mathcal{S}})\leq\gamma(x),\;\forall x\in\mathcal{D}\setminus\mathcal{S},$
which follows directly from the above minimization as $x$ is a feasible point if $\|x\|_{\mathcal{S}}>0$. Additionally, it holds that $p$ is non-decreasing with increasing $r$, since $\tilde{r}>r$ implies $\mathcal{D}\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}<\tilde{r}\}\subseteq\mathcal{D}\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}<r\}$, i.e., the domain in the constrained minimization problem (second case in (12)) is non-increasing. As $p(r)$ is not necessarily continuous for non-convex sets $\mathcal{S}$, we construct a class $\mathcal{K}$ function $\alpha:[0,\overline{r}]\to\mathbb{R}_{\geq 0}$ which is a lower bound of $p$ on $\mathcal{D}\setminus\mathcal{S}$ similarly to the result in [37, Lemma 2.5]. Consider the partitions $(R_{k+1},R_{k}]$ on the interval $(0,\overline{r}]$ with boundary points $R_{k}:=\overline{r}2^{-k},\mathrm{for}\;k=0,1,2,\dots$ and define $\phi_{k}:=2^{-k}p(R_{k+1})$. Note that $\phi_{k+1}<\phi_{k}$, since $p(r)$ is non-decreasing and $2^{-k-1}<2^{-k}$ $\forall k\geq 0$. In addition, it holds that $\phi_{k}\leq p(r),\forall r\in(R_{k+1},R_{k}]$ as $\phi_{k}=2^{-k}p(R_{k+1})\leq p(R_{k+1})$.
We construct the function
$\alpha(r)=\begin{cases}0&\text{if }r=0,\\ \phi_{k+1}+\frac{\phi_{k}-\phi_{k+1}}{(R_{k}-R_{k+1})}(r-R_{k+1}),&r\in(R_{k+1},R_{k}],\end{cases}$
which is piece-wise affine on successive intervals $r\in(R_{k+1},R_{k}]$, continuous, as $\lim_{k\rightarrow\infty}\phi_{k}=0$, strictly increasing and is therefore a class $\mathcal{K}$ function. Furthermore, it holds that $\gamma(r)\leq p(r)$, which implies $\alpha(\|x\|_{\mathcal{S}})\leq\gamma(x)\;\forall x\in\mathcal{D}\setminus\mathcal{S}$ as desired.
In a similar fashion, an upper bound for positive continuous functions can be constructed, which to the best of the author’s knowledge has not been previously proposed in this form. Note that in [8, Proposition B.25], such an upper bound is shown only with respect to a singleton $\mathcal{S}=\{0\}$.
Lemma 4.4.
Consider the compact and non-empty sets $\mathcal{S}$ and $\mathcal{D}$, with $\mathcal{S}\subset\mathcal{D}$ and a continuous function $\gamma:\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}\rightarrow\mathbb{R}_{\geq 0}$, with $\gamma(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$ and $\gamma(x)=0$ for all $\partial_{\mathcal{D}}\mathcal{S}$. Then there exists a $\mathcal{K}$-function $\beta:[0,\bar{r}]\rightarrow\mathbb{R}_{\geq 0}$, with $\bar{r}:=\max_{x\in\mathcal{D}}\|x\|_{\mathcal{S}}$ such that
$\gamma(x)\leq\beta(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}$ (13)
Proof 4.5.
Define the auxiliary function $q\!:\![0,\bar{r}]\rightarrow\mathbb{R}_{\geq 0}$ as
$q(r):=\max\limits_{y}\;\;$ $\gamma(y)$ (14)
s.t. $y\in(\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S})\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}>r\}.$
As $\gamma$ is a continuous function and $(\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S})\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}>r\}$ is a compact set, the above maximum is attained for all $r\in[0,\bar{r}]$. It holds that
$q(\|x\|_{\mathcal{S}})\geq\gamma(x)\;\;\forall x\in\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}$
which follows directly from the fact that $x$ is a feasible point in the above maximisation and the fact that $\gamma(x)=0,\;\forall x\in\partial_{\mathcal{D}}\mathcal{S}$. Additionally, it holds that $q$ is non-decreasing with $r$, since $\tilde{r}>r$ implies $(\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S})\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}>r\}\subseteq(\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S})\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}>\tilde{r}\}$, i.e., the domain in the constrained maximisation problem in (14) is non-decreasing.
A continuous, strictly increasing upper bound for $q(r)$ is then given in the proof of [37, Lemma 2.5] by defining the partitions $[R_{k+1},R_{k})$ with $R_{k}=2^{-k}\bar{r}$ for $k\geq 0$, the quantities $\phi_{k}=q(R_{k})+2^{-k}$ for $k\geq 0$ and $\phi_{-1}=q(\bar{r})+2=\phi_{0}$ and defining the piece-wise affine function
$\beta(r)=\begin{cases}0&\textup{ if }r=0\\ \phi_{k}+\frac{\phi_{k-1}-\phi_{k}}{R_{k}-R_{k+1}}(r-R_{k+1})&\textup{ if }r\in[R_{k+1},R_{k})\\ \phi_{-1}&\textup{ if }r=\bar{r}\end{cases}$ (15)
It then holds that $\beta(0)=0$, $\beta$ is strictly increasing as $\phi_{k-1}\geq\phi_{k}$ and $2^{-k+1}>2^{-k}$ and continuous as $\lim_{r\rightarrow 0^{+}}\beta(r)=0$ as $\lim_{k\rightarrow\infty}\phi_{k}=0$ and $\lim_{r\rightarrow\bar{r}^{-}}\beta(r)=\beta(\bar{r})$, therefore we have $\beta\in\mathcal{K}$.
[…]
Figure 1: Illustrative example of the class $\mathcal{K}$ lower and upper bounds in Lemmata 1 and 4.4. The function $\gamma(x)$ is minimised on $\mathcal{D}\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}<r\}$ (green) and maximised on $(\mathcal{D}\setminus\textup{int}_{\mathcal{D}}\mathcal{S})\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}>r\}$ (grey) to obtain the non-decreasing functions $p(\|x\|_{\mathcal{S}})$ and $q(\|x\|_{\mathcal{S}})$, respectively. Finally, the class $\mathcal{K}$ bounds are constructed via piecewise affine lower and upper bounds.
An illustrative example of the constructed upper and lower bounds is provided in Figure 1. We can now use the results in Lemmata 1 and 4.4, to show that for any function $h$, which is a CBF according to Definition 1, the function $\tilde{h}(x):=\max\{0,h(x)\}$ is a CBF according to Definition 4.
Theorem 4.6.
Consider a CBF $h:\mathcal{D}\rightarrow\mathbb{R}$ according to Definition 1 and a corresponding decrease function $\Delta h$. It holds that the function $\tilde{h}:\mathcal{D}\rightarrow\mathbb{R}_{\geq 0}$ given by $\tilde{h}(x):=\max\{0,h(x)\}$ is a CBF according to Definition 4.
Proof 4.7.
We prove the three statements in Definition 4 for $\tilde{h}(x)=\max\{0,h(x)\}$ based on the fact that $h$ fulfills Definition 1:
(1) From Definition 1, it follows that the safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}=\tilde{\mathcal{S}}=\{x\in\mathbb{R}^{n}\mid\tilde{h}(x)=0\}$ and $\mathcal{D}$ are compact and non-empty.
(2) From Lemma 1, it follows that $\exists\;\alpha_{1}\in\mathcal{K}$, such that $\alpha_{1}(\|x\|_{\mathcal{S}})\leq h(x)=\tilde{h}(x)\;\forall x\in\mathcal{D}\setminus\mathcal{S}$. As $\tilde{h}(x)=0\;\forall x\in\mathcal{S}$, it holds that $\alpha_{1}(\|x\|_{\mathcal{S}})\leq\tilde{h}(x)\;\forall x\in\mathcal{D}$. From Lemma 4.4, it then holds that as $h(x)$ is continuous on $\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}$ and $h(x)=0\;\forall x\in\partial_{\mathcal{D}}\mathcal{S}$ by definition, $\exists\;\alpha_{2}\in\mathcal{K}$ such that $h(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}$. With $\tilde{h}(x)=0\;\forall x\in\mathcal{S}$, it holds that $\tilde{h}(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}$.
(3) We prove a class $\mathcal{K}$ decrease function for $\tilde{h}$ for all $x\in\mathcal{D}$ by considering the following case distinction, where we denote an optimal input as $u^{*}\in\arg\min_{u\in\mathcal{U}}\{h(f(x,u))\mid f(x,u)\in\mathcal{D}\}$:
  • $\forall x\in\mathcal{D}\setminus\mathcal{S}$, we consider the two following cases
    • $f(x,u^{*})\in\mathcal{D}\setminus\mathcal{S}:$
      $\inf_{u\in\mathcal{U}}\{h(f(x,u))\mid f(x,u){\in}\mathcal{D}\}-h(x){\leq}{-}\Delta h(x)$
      ${\Rightarrow}\!\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u)){\mid}f(x,u){\in}\mathcal{D}\}{-}\tilde{h}(x){\leq}{-}\Delta h(x)$. As $\Delta h$ fulfils the conditions of Lemma 1, it holds that $\exists\beta\in\mathcal{K}$, such that $\Delta h(x)\geq\beta(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\mathcal{S}$ and
      $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u))\mid f(x,u){\in}\mathcal{D}\}-\tilde{h}(x)\leq-\beta_{1}(x)$.
    • $f(x,u^{*})\in\mathcal{S}:$ It holds that $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u))\mid f(x,u)\in\mathcal{D}\}=0<\tilde{h}(x)$. Using the lower bound $\alpha_{1}(\|x\|_{\mathcal{S}})\leq\tilde{h}(x)$ from (2), it holds that $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u)){\mid}f(x,u){\in}\mathcal{D}\}{-}\tilde{h}(x){\leq}{-}\alpha_{1}(\|x\|_{\mathcal{S}})$
    We can therefore define the function $\alpha_{3}(\|x\|_{\mathcal{S}})=\min\{\beta(\|x\|_{\mathcal{S}}),\alpha_{1}(\|x\|_{\mathcal{S}})\}\in\mathcal{K}$, such that
    $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u))\mid f(x,u){\in}\mathcal{D}\}-\tilde{h}(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}})$
  • $\forall x\in\mathcal{S}:\;\inf_{u\in\mathcal{U}}h(f(x,u))\leq 0$
    $\Rightarrow\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u)){\mid}f(x,u){\in}\mathcal{D}\}{-}\underbrace{\tilde{h}(x)}_{=0}{\leq}\underbrace{{-}\alpha_{3}(\|x\|_{\mathcal{S}})}_{=0}$ where we used $h(x)\leq\tilde{h}(x)=0\;\forall x\in\mathcal{S}$.
We therefore have $\forall x\in\mathcal{D}$, that
$\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u))\mid f(x,u){\in}\mathcal{D}\}-\tilde{h}(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}})$
Given that we have shown the relation between the two considered definitions for CBFs, it follows that the optimal value function $h_{\mathrm{PB}}$ in (10) is also a CBF according to Definition 4.
Corollary 4.8.
Let Assumptions 1 and 2 hold. The minimum (10) exists and if $\alpha_{f}$ is chosen sufficiently large, then the optimal value function $h_{\mathrm{PB}}(x(k))$ defined in (10a) is a (predictive) control barrier function according to Definition 4 with domain $\mathcal{D}_{\mathrm{PB}}:=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)\leq\alpha_{f}\gamma_{f}\}$ and safe set $\mathcal{S}_{\mathrm{PB}}:=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)=0\}$.
Proof 4.9.
This result follows directly from Theorem 2, which shows that $h_{\mathrm{PB}}$ is a CBF according to Definition 1 under the given conditions. As $h_{\mathrm{PB}}=0$ for all $x\in\mathcal{S}_{\mathrm{PB}}$ and $h_{\mathrm{PB}}>0$ for all $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, it holds that $h_{\mathrm{PB}}(x)=\max\{0,h_{\mathrm{PB}}(x)\}$ and $h_{\mathrm{PB}}$ is a CBF according to Definition 4 by Theorem 4.6.
4.2 Inherent robustness
In this section, we use the relation in Theorem 4.6 to analyse the inherent robustness properties of the PCBF algorithm with respect to external disturbances. With inherent robustness, we denote that a controller, which was designed with nominal system dynamics (1) maintains robust stability guarantees when the closed-loop system is subject to disturbances. Consider therefore the following difference inclusion
$x(k+1)$ $\in F(x(k),w(k))$ (16)
=$=\{x^{+}\in\mathbb{R}^{n}\mid$ $x^{+}=\tilde{f}(x(k),u,w(k))$
$u\in\mathcal{U}_{h_{\mathrm{PB}}}\},$
where $\tilde{f}:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{n}$ is continuous with $\tilde{f}(x,u,0)=f(x,u)$, $w(k)\in\mathbb{R}^{p}$ is an exogeneous disturbance and $\mathcal{U}_{h_{\mathrm{PB}}}$ is the set of safe inputs induced by $h_{\mathrm{PB}}$ according to Definition 2. The aim is to show that if in closed-loop, the inputs $u(k)$ are selected from the set $\mathcal{U}_{h_{\mathrm{PB}}}$, the closed-loop system is input-to-state stable (ISS) with respect to the disturbance $w$. For input-to-state stability, we use the following comparator function definition.
Definition 4.10 (Definition B.42 in [8]).
The system $x(k+1)\in F(x(k),w(k))$ is input-to-state stable in $\mathcal{D}$ with respect to the compact, non-empty and 00-input invariant set [^2] A set $\mathcal{S}\subset\mathcal{D}$ is denoted 00-input invariant for the dynamics $x(k+1)\in F(x(k),w(k))$ if it is invariant for the dynamics $x(k+1)\in F(x(k),0)$. $\mathcal{S}$ if there exists a $\mathcal{KL}$ function [^3] According to [8, Definition B.3], a function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{N}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ belongs to class $\mathcal{KL}$ if it is continuous and if, for each $t\geq 0$, $\beta(\cdot,t)$ is a class $\mathcal{K}$ function and for each $s\geq 0$, $\beta(s,\cdot)$ is non-increasing and satisfies $\lim_{t\rightarrow\infty}\beta(s,t)=0$. $\beta$ and a $\mathcal{K}$ function $\sigma$ such that, for each $x\in\mathcal{D}$, with $\mathcal{D}$ compact and robustly positively invariant, and each disturbance sequence $\mathbf{w}=(w(0),w(1),\dots)$ in $\ell_{\infty}$
$\|\phi(i;x(0),\mathbf{w}_{i})\|_{\mathcal{S}}\leq\beta(\|x\|_{\mathcal{S}},i)+\sigma(\|\mathbf{w}_{i}\|)$
for all $i\in\mathbb{N}_{\geq 0}$, where $\phi(i;x(0),\mathbf{w}_{i})$ is any solution of (16) at time $i$, if the initial state is $x(0)$ and the input sequence is $\mathbf{w}_{i}:=(w(0),w(1),\dots,w(i-1))$.
We note that with regards to input-to-state stability, the disturbances $w(k)$ are considered as the input of the system. Following the definition of ISS, a corresponding ISS Lyapunov function can be defined, which we denote as ISS control barrier function in this paper to distinguish stability of singletons from stability of sets.
Definition 4.11 (Definition B.43 in [8]).
Let $\mathcal{S}$ and $\mathcal{D}$ be compact and non-empty, with $\mathcal{S}$ 00-input invariant, $\mathcal{D}$ robustly positively invariant and $\mathcal{S}\subset\mathcal{D}$. A function $h:\mathcal{D}\rightarrow\mathbb{R}_{\geq 0}$ is an ISS-CBF for system $x(k+1)\in F(x(k),w(k))$ if there exist class $\mathcal{K}$ functions $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ and $\sigma$ such that for all $x\in\mathcal{D},w\in\mathbb{R}^{p}$,
$\alpha_{1}(\|x\|_{\mathcal{S}})$ $\leq h(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}})$ (17a)
$\sup_{x^{+}\in F(x,w)}h(x^{+})$ $-h(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}})+\sigma(\|w\|)$ (17b)
The following Lemma establishes the relationship between the existence of an ISS-CBF and ISS of the system.
Lemma 4.12 (Lemma B.44 in [8]).
Suppose $\tilde{f}$ is continuous and that there exists a continuous ISS-CBF according to Definition 4.11 for $x(k+1)\in F(x(k),w(k))$. Then the system $x(k+1)\in F(x(k),w(k))$ is ISS.
Remark 4.13.
Note that in [8, Appendix B], the definitions of ISS and of an ISS-CBF are given with respect to the singleton $\mathcal{S}=\{0\}$, however this result is applicable also for compact sets $\mathcal{S}$, see, e.g., [38, Remark 3.12]. Additionally, the results are given for the difference equation $x(k+1)=f(x(k),w(k))$ rather than the difference inclusion $x(k+1)\in F(x(k),w(k))$, but this difference is minor due to the supremum in the decrease condition (17b).
By leveraging the fact that the PCBF defined in (10) is a CBF according to Definition 4 according to Corollary 4.8, we are able to relate it to the notion of ISS-CBF in Definition 4.11.
Theorem 4.14.
Let Assumptions 1 and 2 hold. The closed-loop system (16) is ISS according to Definition 4.10.
Proof 4.15.
We show ISS by proving that $h_{\mathrm{PB}}$, as defined in (10), is an ISS-CBF according to Definition 4.11. From Corollary 4.8, it follows that $h_{\mathrm{PB}}$ is a CBF according to Definition 4. It therefore holds that $\mathcal{S}_{\mathrm{PB}}$ is 00-input invariant for the difference inclusion $x(k+1)\in F(x(k),0)$ as it holds that $u\in\mathcal{U}_{h_{\mathrm{PB}}}$ according to (16). Furthermore, from Definition 4, it follows that there exist $\alpha_{1}$ and $\alpha_{2}$ such that $\alpha_{1}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})\leq h_{\mathrm{PB}}(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})$. It therefore remains to be shown that $\sup_{x^{+}\in F(x,w)}h_{\mathrm{PB}}(x^{+})-h_{\mathrm{PB}}(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})+\sigma(\|w\|)$. Given the definition of $\mathcal{U}_{h_{\mathrm{PB}}}$, it holds that $\forall x\in\mathcal{D}_{\mathrm{PB}}$, we have
$\sup_{x^{+}\in F(x,w)}h_{\mathrm{PB}}(x^{+})=\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,w))$
From the Heine-Cantor theorem in [39, Theorem 4.19], it follows that the continuous functions $h_{\mathrm{PB}}$ and $f$ are uniformly continuous on the compact set $\mathcal{D}_{\mathrm{PB}}$. It therefore holds that there exist class $\mathcal{K}$ functions $\sigma_{h_{\mathrm{PB}}}$ and $\sigma_{f}$ according to [40, Lemma 1] such that
$\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,w))$
$\leq$ $\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,0))+\sigma_{h_{\mathrm{PB}}}\circ\sigma_{f}(\|w\|)$
$=$ $\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,0))+\sigma(\|w\|),$
where $\sigma=\sigma_{h_{\mathrm{PB}}}\circ\sigma_{f}$ is also a class $\mathcal{K}$ function. By using $\mathcal{U}_{h_{\mathrm{PB}}}=\{u\in\mathcal{U}\mid h_{\mathrm{PB}}(\tilde{f}(x,u,0))-h_{\mathrm{PB}}(x)\leq-\Delta h_{\mathrm{PB}}(x)\}$ for $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, we have
$\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,0))+\sigma(\|w\|)$
$\leq$ $h_{\mathrm{PB}}(x)-\Delta h_{\mathrm{PB}}(x)+\sigma(\|w\|)$
$\leq$ $h_{\mathrm{PB}}(x)-\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})+\sigma(\|w\|),$
where $\alpha_{3}$ exists according to Lemma 1. For any $x\in\mathcal{S}_{\mathrm{PB}}$, we have $\mathcal{U}_{h_{\mathrm{PB}}}=\{u\in\mathcal{U}\mid h_{\mathrm{PB}}(\tilde{f}(x,u,0))\leq 0\}$, resulting in
$\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,0))+\sigma(\|w\|)$
$\leq\sigma(\|w\|)$
$=h_{\mathrm{PB}}(x)-\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})+\sigma(\|w\|),$
as $h_{\mathrm{PB}}(x)=\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})=0$ for $x\in\mathcal{S}_{\mathrm{PB}}$. Therefore, for any $x\in\mathcal{D}_{\mathrm{PB}}$ the decrease condition (17b) holds and $h_{\mathrm{PB}}$ is an ISS-CBF according to Definition 4.11 and the closed-loop system (16) is ISS, which follows from Lemma 4.12.
Given that the PCBF defines an ISS-CBF for system (16), it holds that the nominal predictive control problem (10) admits inherent robustness properties. In other words, it holds that for any state $x(0)\in\mathcal{D}_{\mathrm{PB}}$, a neighbourhood of $\mathcal{S}_{\mathrm{PB}}$ is stable for the system (16).
5 EXPLICIT APPROXIMATION OF THE PCBF
In this section, we present the main proposed approximation-based algorithm. Given the theoretical foundation established in Section 4, it holds that a closed-loop system (16) is inherently robust with respect to exogeneous disturbances and in this section, we show how possible approximation errors can be treated in a similar manner. Furthermore, we propose a continuous extension for the PCBF, which is negative whenever $h_{\mathrm{PB}}$ is 00, such that safety filter interventions due to approximation errors within $\mathcal{S}_{\mathrm{PB}}$ are reduced.
5.1 Approximate PCBF
The optimal value function in (10) is a CBF according to Definition 4 and encodes the information of the entire prediction horizon of the predictive optimisation problem with respect to potential constraint violations. Therefore, a time horizon independent optimization problem that could be solved instead of (10) is given by
$\min_{u\in\mathcal{U}}h_{\mathrm{PB}}(f(x(k),u))$ (18)
Clearly, such a formulation benefits from the same guarantees as solving (10). Given a state $x\in\mathcal{S}_{\mathrm{PB}}$, any input $u\in\arg\min h_{\mathrm{PB}}(f(x(k),u))$ ensures that at the next time step, the system remains in the safe set $\mathcal{S}_{\mathrm{PB}}$, rendering it invariant. Similarly, for states in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, by minimising the PCBF value, the decrease $\Delta h_{\mathrm{PB}}$ in (3a) is achieved and stability of $\mathcal{S}_{\mathrm{PB}}$ is ensured.
However, an explicit formulation of $h_{\mathrm{PB}}$ is generally not available. We therefore propose to approximate the function $h_{\mathrm{PB}}$ with a continuous regressor $\hat{h}_{\mathrm{PB}}$. Due to the continuity property of the PCBF in (10) and the fact that it can be sampled offline, the function is suitable for approximation, such that a computationally efficient algorithm can be leveraged. As the approximation $\hat{h}_{\mathrm{PB}}$ will be embedded in an optimisation problem, we consider neural networks in this paper due to the parametric nature of the regressor, allowing for a formulation independent of the number of training samples, compared to, e.g., kernel-based methods. Note that while in principle a safety filter policy $\pi(x,u_{\mathrm{p}})$ could be learned, this policy is generally not only non-continuous and multi-valued, but also has a higher input dimensionality for the data collection and training as $\pi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$. As an illustrative example, consider a car approaching an obstacle. If the performance input is given to steer straight towards the obstacle, an optimal solution for the safety filter problem is given by either steering fully to the left or to the right.
In order to train the regressor for the PCBF approximation, a dataset of samples has to be collected offline in the form of $\mathbb{D}=\{(x_{i},h_{\mathrm{PB}}(x_{i}))\}_{i=0}^{n_{d}}$. This data set can be sampled from the entire region of attraction $\mathcal{D}_{\mathrm{PB}}=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)\leq\gamma_{f}\alpha_{f}\}$, where $\gamma_{f}$ is the sublevel set of the domain of the terminal CBF $\mathcal{D}_{f}$ and $\alpha_{f}$ is the terminal penalty in (10a), or from a subset deemed relevant to the specific application. In the Appendix, we propose a geometric sampling method based on [41], which allows efficiently sampling the PCBF by performing a pseudo random walk over a preselected domain, using previous solutions as warmstarts to the next computed optimisation problem.
Given a trained explicit approximation $\hat{h}_{\mathrm{PB}}$ of $h_{\mathrm{PB}}$, the optimisation problem (18) could directly be modified to the optimisation problem $\min_{u\in\mathcal{U}}\hat{h}_{\mathrm{PB}}(f(x(k),u))$. Such an approach can be useful as an efficient control mechanism for convergence of states $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ to $\mathcal{S}_{\mathrm{PB}}$ with no approximation errors, however, it does not easily allow for considering the primary objective considered in this paper, i.e. filtering inputs with respect to safety regards. Solving a subsequent safety filter problem (9) would still require the full information of the optimal slack variables across the prediction horizon. Instead of minimising the approximate PCBF value, the corresponding CBF decrease condition (3a) can directly be used as a constraint as is standard in CBF-based safety filters, see, e.g. [5] for continuous-time CBF applications. The optimisation problem which is solved at every time step is then given by
$\begin{split}\min_{u}&\|u-u_{\mathrm{p}}(k)\|\\ \textup{s.t. }&\hat{h}_{\mathrm{PB}}(f(x(k),u))\!-\!\hat{h}_{\mathrm{PB}}(x(k))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x(k)),\\ &u\in\mathcal{U},\end{split}$ (19)
where $c_{\Delta}\in(0,1]$ is a user defined hyperparameter and the function $\Delta\hat{h}_{\mathrm{PB}}(x)$ is a decrease function of the approximate PCBF $\hat{h}_{\mathrm{PB}}(x)$. The main advantage of the proposed optimisation problem compared to the original PCBF mechanism, which solves (10) and (9) consecutively consists of the fact that the formulation in (19) is independent of the prediction horizon of the predictive control based optimisation problems. Instead, only a single decrease constraint needs to be fulfilled along with the input constraints and the only optimisation variable is $u\in\mathbb{R}^{m}$.
Given that the decrease function of $h_{\mathrm{PB}}(x)$, i.e. $\Delta h_{\mathrm{PB}}(x)$, is not explicitly available in a similar fashion to $h_{\mathrm{PB}}(x)$ itself, we propose an optimisation-based formulation for $\Delta\hat{h}_{\mathrm{PB}}$. The approximate decrease which is used in (19) is defined as
$\Delta\hat{h}_{\mathrm{PB}}(x):=\max_{u\in\mathcal{U}}\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u))$ (20)
From this formulation of the approximate decrease function, it follows that the hyperparameter $c_{\Delta}$ allows tuning the desired decrease rate of the closed-loop system for any $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$. By selecting $c_{\Delta}=1$, at every time step, the approximate PCBF $\hat{h}_{\mathrm{PB}}$ undergoes the maximal possible decrease, while choosing $c_{\Delta}\in(0,1)$ reduces the safety interventions of the proposed safety filter by reducing the amount that the approximate PCBF is required to decrease. This provides an additional degree of freedom compared to the PCBF mechanism, where at every time steps the minimal slacks in terms of the objective function (10a) are required to be used in the safety filter and relaxing the corresponding decrease is a theoretically challenging problem.
Remark 5.16.
In principle, a decrease function $\Delta\hat{h}_{\mathrm{PB}}(x)$ can also be hand-tuned as is common in the continuous-time CBF literature, such as, e.g. choosing an exponential decrease $\Delta\hat{h}_{\mathrm{PB}}(x)=\alpha(h_{\mathrm{PB}}(x))$ for some class $\mathcal{K}$ function $\alpha$, see, e.g., [5]. Such a hand-tuned decrease function reduces the computational demand of the proposed scheme by only requiring the solution to one optimisation problem. However, in this case recursive feasibility, which will be shown later in this section for the proposed maximal decrease can no longer be ensured.
An additional advantage of the explicit approximation of the PCBF is given by the fact that proposed inputs $u_{\mathrm{p}}(k)$ can be very efficiently certified in terms of satisfying optimality with respect to (19) if $u_{\mathrm{p}}(k)\in\mathcal{U}$. It suffices to perform a simple evaluation, whether $\hat{h}_{\mathrm{PB}}(f(x(k),u_{\mathrm{p}}(k)))\leq 0$, as it then holds that $u_{\mathrm{p}}(k)$ is an optimal solution to (19). This function evaluation can thereby reduce the need for solving two optimisation problems online at every time step. The main proposed algorithm is then given in Algorithm 1, where first, the proposed input is certified and if the input is not deemed to be safe, the maximal decrease in (20) is computed and then used in the safety filter (19).
Algorithm 1 Approximate PCBF safety filter
1:  for $k\geq 0$ do
2:     Measure state $x(k)$
3:     Obtain the proposed control input $u_{\mathrm{p}}(k)$
4:     if $u_{\mathrm{p}}(k)\in\mathcal{U}\wedge\hat{h}_{\mathrm{PB}}(f(x(k),u_{\mathrm{p}}(k)))\leq 0$ then
5:        Apply $u(k)\leftarrow u_{\mathrm{p}}(k)$
6:     else
7:        Compute $\Delta\hat{h}_{\mathrm{PB}}(x(k))$ using (20)
8:        Apply $u(k)\leftarrow u^{*}$, solution of (19)
9:     end if
10:  end for
Given the proposed algorithm, we can now provide theoretical guarantees of the closed-loop behaviour of the system despite possible approximation errors. First, we show that the optimisation problem (19) is feasible using the maximum decrease (20) for any $x\in\mathcal{D}_{\mathrm{PB}}$.
Proposition 5.17.
The optimisation problem (19), where $\Delta\hat{h}_{\mathrm{PB}}(x)$ is defined in (20) is feasible for any $x(k)\in\mathbb{R}^{n}$ and any $c_{\Delta}\in(0,1]$.
Proof 5.18.
The optimisation problem (20) is feasible for any $x\in\mathbb{R}^{n}$ by choosing any $u\in\mathcal{U}$. Given an optimal solution $u^{*}$ to (20), it holds that $u^{*}$ is feasible in (19) as $u^{*}\in\mathcal{U}$ and for any $x\in\mathbb{R}^{n}$, $\hat{h}_{\mathrm{PB}}(f(x,u^{*}))-\hat{h}_{\mathrm{PB}}(x)=-\max_{u\in\mathcal{U}}\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x)$ for all $c_{\Delta}\in(0,1]$.
In a second step, we can show that the closed-loop system is ISS with respect to the approximation errors of $\hat{h}_{\mathrm{PB}}$ if the approximation error is bounded according to the following assumption.
Assumption 3
The approximation error of $\hat{h}_{\mathrm{PB}}$ is uniformly bounded, i.e., it holds that $|h_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(x)|\leq\epsilon_{h}$ for all $x\in\mathcal{D}_{\mathrm{PB}}$. In other words, it holds that $\hat{h}_{\mathrm{PB}}(x)=h_{\mathrm{PB}}(x)+e_{h}(x)$, where the error $e_{h}(x)$ is bounded by $|e_{h}(x)|\leq\epsilon_{h}$.
Such a uniform bound for the approximation error always holds if the chosen regressor for $\hat{h}_{\mathrm{PB}}$ is continuous. Due to continuity of $\hat{h}_{\mathrm{PB}}$ and $h_{\mathrm{PB}}$ as well as compactness of $\mathcal{D}_{\mathrm{PB}}$, a maximum $\epsilon_{h}$ exists for $\max_{x\in\mathcal{D}_{\mathrm{PB}}}|h_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(x)|$ according to the extreme value theorem [8, Proposition A.7].
Theorem 5.19.
Let Assumptions 12 and 3 hold. Then, for small enough approximation error $\epsilon_{h}$, the closed-loop system (1) under application of Algorithm 1 is ISS according to Definition 4.10.
Proof 5.20.
We show that $h_{\mathrm{PB}}(x)$ is an ISS Lyapunov function according to Definition 4.11 for the closed-loop system $x(k+1)\in\hat{F}(x(k))=\{x^{+}\in\mathbb{R}^{n}\mid x^{+}=f(x(k),u),u\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}\}$, where $\mathcal{U}_{\hat{h}_{\mathrm{PB}}}\coloneqq\{u\in\mathcal{U}\mid\hat{h}_{\mathrm{PB}}(f(x(k),u))\!-\!\hat{h}_{\mathrm{PB}}(x(k))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x(k))\}$. Given that inputs are selected as an optimal solution to (19), it holds that $u\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}$ under application of Algorithm 1. As $h_{\mathrm{PB}}$ is a CBF according to Definition 4, the condition (17a) is satisfied. Furthermore, we can show that for any $u\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}$, it holds that
$h_{\mathrm{PB}}(f(x,u))-h_{\mathrm{PB}}(x)$
$\leq$ $\hat{h}_{\mathrm{PB}}(f(x,u))-e_{h}(f(x,u))-\hat{h}_{\mathrm{PB}}(x)+e_{h}(x)$
$\leq$ $-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x)+2\epsilon_{h}$
$\leq$ $-c_{\Delta}\max_{u\in\mathcal{U}}\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u))+2\epsilon_{h}$
$\leq$ $-c_{\Delta}\Delta h(x)+4\epsilon_{h},$
where we used $c_{\Delta}\in(0,1]$ in the last inequality, such that (17b) is satisfied. The set $\mathcal{S}_{\mathrm{PB}}$ is 00-invariant with $\epsilon_{h}=0$ and the set $\mathcal{D}_{\mathrm{PB}}$ is robust invariant, if $\epsilon_{h}$ is sufficiently small, i.e., for all $x\in\mathcal{D}_{\mathrm{PB}}=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)\leq\alpha_{f}\gamma_{f}\}$, it holds that $h_{\mathrm{PB}}(f(x,u))\leq\alpha_{f}\gamma_{f}$. This holds if $\epsilon_{h}\leq\frac{1}{4}(\alpha_{f}\gamma_{f}-\bar{c})$ with $\bar{c}\coloneqq\max_{x\in\mathcal{D}_{\mathrm{PB}}}h_{\mathrm{PB}}(x)-c_{\Delta}\alpha_{3}\circ\alpha_{1}^{-1}(h_{\mathrm{PB}}(x))\leq\alpha_{f}\gamma_{f}$ as we have $h_{\mathrm{PB}}(x)\leq\alpha_{f}\gamma_{f}$ and
$h_{\mathrm{PB}}(x)-c_{\Delta}\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})+4\epsilon_{h}$
$\leq$ $h_{\mathrm{PB}}(x)-c_{\Delta}\alpha_{3}\circ\alpha_{1}^{-1}(h_{\mathrm{PB}}(x))+4\epsilon_{h}\stackrel{{\scriptstyle!}}{{\leq}}\alpha_{f}\gamma_{f}$
It therefore holds that using the approximation-based safety filter (19), the system converges to and remains in a neighbourhood of the safe set $\mathcal{S}_{\mathrm{PB}}$.
Remark 5.21.
In the proof of Theorem 5.19, the error $|e_{h}(x)|$ can also be bounded by a class $\mathcal{K}$ function $\gamma$ and a constant term $\tilde{\epsilon}_{h}\leq\epsilon_{h}$, such that $|e_{h}(x)|\leq\gamma(x)+\tilde{\epsilon}_{h}$. If it holds that $\forall x\in\mathcal{D}_{\mathrm{PB}}$, $c_{\Delta}\Delta h_{\mathrm{PB}}(x)>2\gamma(x)$, the decrease condition is still satisfied with respect to the smaller constant $2\tilde{\epsilon}_{h}+2\epsilon_{h}$. This implies that approximation errors close to $\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$ have a bigger impact on the closed-loop stability than errors when $h_{\mathrm{PB}}$ is large, which can be leveraged when performing the approximation as discussed in Section 6. Furthermore, it can be shown that the approximation is additionally input to state stable with respect to an additional disturbance $w$ acting on system (1), following the results in Theorem 4.14.
5.2 Continuous extension of the PCBF
As the PCBF $h_{\mathrm{PB}}(x)=0$ for all $x\in\mathcal{S}_{\mathrm{PB}}$, errors in its approximation can lead to undesired conservativeness of the safety filter within $\mathcal{S}_{\mathrm{PB}}$ if $\hat{h}_{\mathrm{PB}}(x)>0$, as potentially safe control inputs can be deemed unsafe. In order to avoid such unnecessary safety filter interventions, we propose a continuous extension of $h_{\mathrm{PB}}$ in $\mathcal{S}_{\mathrm{PB}}$, which is negative and therefore mitigates potential learning errors leading to the state being considered unsafe. Consider the function
$h(x):=\begin{cases}h_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}\\ \bar{h}_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{S}_{\mathrm{PB}},\end{cases}$ (21)
where the extension on $\mathcal{S}_{\mathrm{PB}}$ is given by
$\bar{h}_{\mathrm{PB}}(x(k)):=\min_{u_{i|k},\xi_{i|k}}\;\;$ $\max_{i,j}\{[\xi_{i|k}]_{j},\xi_{N|k}\}$ (22a)
s.t. $\forall\,i=0,\dots,N-1,$
$x_{0|k}=x(k),$ (22b)
$x_{i+1|k}=f(x_{i|k},u_{i|k}),$ (22c)
$u_{i|k}\in\mathcal{U},$ (22d)
$x_{i|k}\in\overline{\mathcal{X}}_{i}(\xi_{i|k}),$ (22e)
$h_{f}(x_{N|k})\leq\xi_{N|k}.$ (22f)
Note that the slacks in (22e) and (22f) are not required to be positive, implying negativity of $\bar{h}_{\textup{PB}}$ when the state and terminal constraints are satisfied. We show that the function $h$ is continuous on $\mathcal{D}_{\mathrm{PB}}$ and a control barrier function according to Definition 1.
Theorem 5.22.
Consider the function $h:\mathcal{D}_{\mathrm{PB}}\rightarrow\mathbb{R}$, with
$h(x)=\begin{cases}h_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}},\\ \bar{h}_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{S}_{\mathrm{PB}}.\end{cases}$
It holds that $h$ is a CBF according to Definition 1 and the function $\tilde{h}(x)=\max\{0,h(x)\}=h_{\mathrm{PB}}(x)$ is a CBF according to Definition 4.
Proof 5.23.
We prove continuity of $h$ by proving continuity of $h_{\mathrm{PB}}$ on $\mathcal{D}_{\mathrm{PB}}\setminus\operatorname{int}_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$, continuity of $\bar{h}_{\mathrm{PB}}$ on $\mathcal{S}_{\mathrm{PB}}$ and that $h_{\mathrm{PB}}(x)=\bar{h}_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$.
1) Continuity of $h_{\mathrm{PB}}$ for all $x\in\mathcal{D}_{\mathrm{PB}}\setminus\operatorname{int}_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$: In [12, Theorem III.6 b)], it is shown that $h_{\mathrm{PB}}$ is continuous for all $x\in\mathcal{D}_{\mathrm{PB}}$, which implies continuity of $h$ for all $x\in\mathcal{D}_{\mathrm{PB}}\setminus\operatorname{int}_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$.
2) Continuity of $\bar{h}_{\mathrm{PB}}$ for all $x\in\mathcal{S}_{\mathrm{PB}}$: We can show continuity of $\bar{h}_{\mathrm{PB}}$ by using similar arguments as in [12, Theorem III.6 b)]. For any $\epsilon>0$ such that $|\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})|<\epsilon$, with $x,\bar{x}\in\mathcal{S}_{\mathrm{PB}}$, we show that there exists a $\delta>0$, such that $\|x-\bar{x}\|<\delta$. Consider the optimal solution of (22) at $x$, i.e. $u_{i|k}^{*}(x)$ as a candidate solution in $\bar{x}$, i.e. $\bar{u}_{i|k}(\bar{x})=u^{*}_{i|k}$, which results in the state predictions $\bar{x}_{i|k}(\bar{x})$ with $\bar{x}_{0|k}=\bar{x}$. The resulting slacks are then given by $\bar{\xi}_{i|k}(\bar{x})=\min\{0,c(\bar{x}_{i|k})+\Delta_{i}\mathbf{1}\}$ and $\bar{\xi}_{N|k}=\min\{0,h_{f}(\bar{x}_{N|k}(\bar{x}))\}$. As the initial state and the corresponding inputs lie in a compact set, i.e. $\bar{x}\in\mathcal{S}_{\mathrm{PB}}$ and $\bar{u}_{i|k}(\bar{x})\in\mathcal{U}$, and the functions $f,c_{j}$ and $h_{f}$ are continuous, it holds that the candidate solution also lies within a compact set from [12, Lemma C.3] and from the Heine-Cantor theorem, it follows that $f,c_{j}$ and $h_{f}$ are uniformly continuous. As compositions of functions, the minimum and additions of uniform continuous functions all preserve uniform continuity, the state predictions, the candidate slack variables and the corresponding constructed candidate value function $\tilde{h}_{\mathrm{PB}}(\bar{x})=\max_{i,j}\{[\bar{\xi}_{i|k}(\bar{x})]_{j},\bar{\xi}_{N|k}(\bar{x})\}$ are uniformly continuous in the initial condition $\bar{x}$. From the fact that the optimal solution $\bar{h}_{\mathrm{PB}}(\bar{x})\leq\tilde{h}_{\mathrm{PB}}(\bar{x})$ due to the minimisation in (22), it holds that for any $\epsilon>0$, there exists a uniform $\delta>0$ with $\|x-\bar{x}\|<\delta$, such that $\bar{h}_{\mathrm{PB}}(\bar{x})-\bar{h}_{\mathrm{PB}}(x)\leq\tilde{h}_{\mathrm{PB}}(\bar{x})-\bar{h}_{\mathrm{PB}}(x)<\epsilon$. The same arguments hold when using the optimal solution at $\bar{x}$ in $x$, i.e. $\bar{u}_{i|k}(x)=u_{i|k}^{*}(\bar{x})$, such that there exists $\delta>0$ with $\|x-\bar{x}\|<\delta$, which implies that $\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})\leq\tilde{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})<\epsilon$, which in turn implies that for any $\epsilon>0$, there exists $\delta>0$ with $\|x-\bar{x}\|<\delta$, which implies that $|\bar{h}_{\mathrm{PB}}(x)-\bar{h}_{\mathrm{PB}}(\bar{x})|<\epsilon$.
3) Equality $\bar{h}_{\mathrm{PB}}(x)=h_{\mathrm{PB}}(x)$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$: By definition, we have $h_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$. In order to show that $\bar{h}_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$ we first consider an optimal input sequence $\tilde{u}_{i|k}^{*}(\tilde{x})$ of (22) for any $\tilde{x}\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$. Given this optimal input sequence, we showed in 2)2)2 ) that the corresponding slack variables $\tilde{\xi}_{i|k}(\tilde{x})$ and $\tilde{\xi}_{N|k}(\tilde{x})$ and the optimal value function $\bar{h}_{\mathrm{PB}}(\tilde{x})$ are uniformly continuous in $\tilde{x}$. Furthermore, it holds that the candidate slacks $\tilde{\xi}_{i|k}(\tilde{x})\leq 0$ and $\tilde{\xi}_{N|k}(\tilde{x})\leq 0$ as any optimal inputs $u^{*}_{i|k}(\tilde{x})$ of (10) are feasible in (22) and $h_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}\subseteq\mathcal{S}_{\mathrm{PB}}$. As the candidate slacks are all non-positive, it follows that $\bar{h}_{\mathrm{PB}}(x)\leq 0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$. We can then show that $\bar{h}_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}\subseteq\mathcal{S}_{\mathrm{PB}}$ by contradiction. Suppose that for any $\tilde{x}\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$, it holds that there exists a solution with all slack variables $\xi_{i|k}(\tilde{x})<0$ and $\xi_{N|k}(\tilde{x})<0$ in (22). Then for the corresponding optimal input sequence $\tilde{u}_{i|k}(\tilde{x})$, it holds that $\xi_{i|k}(\tilde{x})$ and $\xi_{N|k}(\tilde{x})$ are continuous in $\tilde{x}$. The continuity of the slack variables in $\tilde{x}$ implies that there exists an open ball of radius $\delta>0$, i.e. $B_{\mathcal{D}}^{\delta}(\tilde{x})=\{x\in\mathcal{D}\mid\|x-\tilde{x}\|<\delta\}$, for which it still holds that $\xi_{i|k}(\tilde{x})<0$ and $\xi_{N|k}(\tilde{x})<0$. As the optimal input $\tilde{u}_{i|k}(\tilde{x})$ is a feasible input for (10), it implies that $h_{\mathrm{PB}}(x)=0$ for all $x\in B_{\mathcal{D}}^{\delta}(\tilde{x})$, which implies that $B_{\mathcal{D}}^{\delta}(\tilde{x})\subseteq\mathcal{S}_{\mathrm{PB}}$. This fact however contradicts with the initial assumption that $\tilde{x}\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$, as by definition any open neighbourhood of $\tilde{x}\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$ must contain at least one point in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$. By contradiction, it therefore holds that any optimal solution of (22) has at least one slack $[\tilde{\xi}^{*}_{i|k}(\tilde{x})]_{j}$ or $\tilde{\xi}^{*}_{N|k}(\tilde{x})$ which is 0 and all slacks are non-positive, which implies that $\bar{h}_{\mathrm{PB}}(\tilde{x})=\max_{i,j}\{[\tilde{\xi}^{*}_{i|k}(\tilde{x})]_{j},\tilde{\xi}^{*}_{N|k}(\tilde{x})\}=0$.
As the function $h$ is a piece-wise function which is continuous on both of its subdomains $\mathcal{S}_{\mathrm{PB}}$ and $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ and $\bar{h}_{\mathrm{PB}}(x)=h_{\mathrm{PB}}(x)=0$ for all $x\in\partial_{\mathcal{D}_{\mathrm{PB}}}\mathcal{S}_{\mathrm{PB}}$, it holds that $h$ is continuous on $\mathcal{D}_{\mathrm{PB}}$.
Remark 5.24.
Note that the function
$h(x):=\begin{cases}h_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}},\\ c_{h}\bar{h}_{\mathrm{PB}}(x)&\textup{ if }x\in\mathcal{S}_{\mathrm{PB}}.\end{cases}$ (23)
is also a CBF according to Definition 1 for any constant $c_{h}\geq 0$ as continuity of $h$ is preserved. The design parameter $c_{h}$ allows for trading off a reduced impact of learning errors in $\mathcal{S}_{\mathrm{PB}}$ on the safety filter interventions and ease of approximation.
In principle, this continuous extension requires solving two optimisation problems for each sample $x\in\mathcal{S}_{\mathrm{PB}}$, i.e., (10) to determine if $x\in\mathcal{S}_{\mathrm{PB}}$ or $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ and (22) in case $x\in\mathcal{S}_{\mathrm{PB}}$. However, a geometric sampler such as used in [41] can be used, which performs a line search over given seed and target points. By using such a geometric sampler, information of the previous solution can be used to solve either (10) or (22) based on which problem was previously solved. It follows that two optimisation problems only need to be solved if the samples on the line cross over from $\mathcal{S}_{\mathrm{PB}}$ into $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ or vice-versa. Additionally, this geometric sampler benefits from being able to warmstart the solver using the previous optimal solution $\mathbf{o}=\{x^{*}_{i},u^{*}_{i},\xi^{*}_{i},\bm{\nu}^{*},\bm{\lambda}^{*}\}$, where $\bm{\nu}^{*}$ and $\bm{\lambda}^{*}$ denote the optimal dual variables. Given that the optimal solution is shown to be continuous in $x$ in [12, Theorem III.6] and Theorem 5.22, the previously computed optimal solution will be close to the solution of the next iteration in the line search. This allows the computation time during sampling to be significantly reduced as shown in [41]. Finally, the random walk can be parallelised by performing line search for different goal points given a random starting seed. The resulting sampling algorithm is then given by Algorithm 2 in the Appendix.
6 NUMERICAL EXAMPLES
6.1 Linear system
As an illustrative example, we consider an unstable linearised pendulum with a damper component
$x(k+1)$ $=Ax(k)+Bu(k)$ (24)
$=\begin{bmatrix}1&T_{s}\\ T_{s}\frac{g}{l}&1-T_{s}c\end{bmatrix}x(k)+\begin{bmatrix}0\\ 5T_{s}\end{bmatrix}u(k),$
where the states are given by the angle and angular velocity, i.e., $x=\left[\psi\;\;\omega\right]^{\top}\in\mathbb{R}^{2}$, the input $u\in\mathbb{R}$ and the dynamics are defined through the sampling time $T_{s}=0.1$, the damping coefficient is $c=1$, the length $l=1.3$ and $g=9.81$. The system is subject to polytopic state and input constraints $x(k)\in\mathcal{X}=\{x\in\mathbb{R}^{2}\mid A_{x}x\leq b_{x}\}$ and $u(k)\in\mathcal{U}=\{u\in\mathbb{R}\mid A_{u}x\leq b_{u}\}$, which are given as bounds on the angle $-0.5\leq\psi\leq 0.5$, the angular velocity $-1.5\leq\omega\leq 1.5$ and the input $-3\leq u\leq 3$. We choose the terminal penalty $\alpha_{f}=1\textup{e}6$ and as tightening values $\Delta_{i}=i\cdot 10^{-3}$ for $i=1,\dots,N$, where the prediction horizon of the PCBF problems (10) and (9) is $N=10$.
PCBF [12]
$\hat{h}_{\mathrm{PB}}$
$c_{\Delta}{=}1$
$\hat{h}$
$c_{\Delta}{=}1$
$\hat{h}_{\mathrm{PB}}$
$c_{\Delta}{=}0.2$
$\hat{h}$
$c_{\Delta}{=}0.2$
$\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ 1 1.05 1.10 0.41 0.48
$\mathcal{S}_{\mathrm{PB}}$ 1 9.33 3.50 1.92 1.35
Table 1: Normalised magnitude of safety interventions for safety filters based on the pcbf and two underfit neural network approximations for 10000 random states
In order to compute the terminal CBF $h_{f}$, we follow a similar procedure to [12, Section IV] by first computing a quadratic Lyapunov function $V(x)=x^{\top}Px$ and corresponding state feedback $u(k)=Kx(k)$ for the system (24). The Lyapunov function is computed such that the ellipse described by $V(x)\leq 1$ is the largest ellipse within the tightened state constraints $\bar{X}_{N-1}(0)$ such that $Kx\in\mathcal{U}$ according to the following semi-definite program
$\min_{E,Y}$ $-\log\det(E)$ (25a)
s.t. $E\succeq 0$ (25b)
$\begin{bmatrix}E&(AE+BY)^{\top}\\ AE+BY&E\end{bmatrix}\succeq 0$ (25c)
$\begin{bmatrix}([b_{x}]_{i}-\Delta_{N-1})^{2}&[A_{x}]_{i}E\\ ([A_{x}]_{i}E)^{\top}&E\end{bmatrix}\succeq 0\;\;\forall i=1,\dots 4$ (25d)
$\begin{bmatrix}[b_{u}]_{j}^{2}&[A_{u}]_{j}Y\\ ([A_{u}]_{j}Y)^{\top}&E\end{bmatrix}\succeq 0\;\;\forall j=1,\dots 2,$ (25e)
similarly to, e.g., [12, Section IV], such that $P$ and $K$ can by computed as $P=E^{-1}$ and $K=YP$. The terminal CBF is then given by $h_{f}=V(x)-1$, such that $\mathcal{S}_{f}$ fulfils Assumption 2 and the domain $\mathcal{D}_{f}$ is given as the largest sublevel set $h_{f}(x)\leq\gamma_{f}\approx 2.95$ for which the input constraints $u=Kx\in\mathcal{U}$ are satisfied for all $x\in\mathcal{D}_{f}$.
[…]
Finally, we compare the closed-loop performance of both algorithms in Figure 2. For this comparison, we train another neural network using the same data sets generated by sampling $h(x)$ by gridding the state space, where we choose a higher expressiveness of the model by using $3$ hidden layers with $16$ softplus activation functions. The model is trained in $30$ epochs using a batch size of $8$, a learning rate of $5\textup{e}-3$ and the MSE loss function, with a total computation time of approximately $25$min. The final validation MSE is given by $6\textup{e}-3$, which is two orders of magnitude lower compared to the two underfit models. As a performance input, we choose $u_{\mathrm{p}}(k)=2$, which aims to drive the system into the constraint $\psi=0.5$ and causes a slight chattering in the velocity at the constraint in closed-loop. By setting $c_{\Delta}=0.15$, we observe that in closed-loop, for 5 different initial conditions using the same initial angle $\psi=0$ and different velocities $\omega$, the system is allowed to converge much slower to $\mathcal{S}_{\mathrm{PB}}$, while still converging to the constraint $\psi=0.5$. The average solve times for the two methods over the $150$ simulated time steps are given by $0.021s$ for the PCBF algorithm in [12], which is solved using CVXPY [43] and MOSEK [44], and $0.011s$ for the approximate PCBF safety filter, which is solved using CasADi [45] and IPOPT [46]. We note that even for the small time horizon of $N=10$, we observe an almost $2$-times computational speedup. While the computation time of the PCBF algorithm is expected to increase with increasing time horizon, the online computation of the approximation based safety filter is independent of the considered time horizon and only the offline data collection time will be impacted.
Figure 2: Closed-loop comparison of the PCBF-based safety filter in (10) and (9) with the proposed approximation scheme for system (24) with states given by the angle $\psi$ (top) and the angular velocity $\omega$ (middle). The explicit approximation enables tuning the rate of the PCBF decrease through the tuning parameter $c_{\Delta}\in(0,1]$, resulting in a slower convergence of the states to $\mathcal{S}_{\mathrm{PB}}$, seen by the corresponding CBF value $h(x)$ (bottom).
6.2 Nonlinear system
In this section, we demonstrate the potential for the proposed approximation-based algorithm as a computationally efficient safety filter. We consider a kinematic miniature race car model discretised using Euler forward and inspired by [47] of the form
$x(k+1)$ $=f(x(k),u(k))$
$=\begin{bmatrix}p_{x}(k)+T_{s}(v(k)\cos(\psi(k)))\\ p_{y}(k)+T_{s}(v(k)\sin(\psi(k)))\\ \psi(k)+T_{s}(v(k)/l\tan(\delta(k)))\\ v(k)+T_{s}u_{0}(k)\\ \delta(k)+T_{s}u_{1}(k)\end{bmatrix},$ (26)
where the length of the car is $l=0.09$m and the sampling time is $T_{s}=0.02$s. The states are given by the $x$- and $y$-position of the car, the heading angle $\psi$, the velocity $v$ and the steering angle $\delta$, i.e. $x=\left[p_{x},p_{y},\psi,v,\delta\right]^{\top}\in\mathbb{R}^{5}$ and the inputs are given by an acceleration and the steering change, resulting in $u\in\mathbb{R}^{2}$. The input constraints are given by $-5\leq u_{0}\leq 2$ and $-2.8\leq u_{1}\leq 2.8$ and the steering lies within $-0.75\leq\delta\leq 0.75$. The car operates in a cluttered environment with 9 obstacles, which are defined by $(p_{x}-o_{x,i})^{2}+(p_{y}-o_{y,i})^{2}\geq r_{i}^{2}$ for each obstacle $i$ with varying centre position $(o_{x,i},o_{y,i})$ and radius $r_{i}$. Finally, there are no constraints on the other states. For compactness of $\mathcal{D}$ and $\mathcal{S}$, they can be assumed to be large enough that they remain inactive during operation.
We consider an increasing tightening $\Delta_{j}=j\cdot 1\textup{e}-4$ over a prediction horizon of $N=50$. As a terminal constraint, we impose the car to stop within the constraints $c_{x}(x)-\Delta_{N}\leq 0$ described above with a velocity $v=0$ and these terminal constraints are not softened with a terminal slack, i.e. $\xi_{N}=0$.
[…]
In order to demonstrate the computational advantages as a safety filter of the proposed approach, we use a naive model predictive control implementation as a performance controller, which is given as
$u_{\mathrm{p}}(k)\in\arg\min_{u_{i|k}}$ $\sum_{i=0}^{30}\|p_{x,i|k}+p_{x}(0)\|^{2}_{2}+\|p_{y,i|k}+p_{y}(0)\|^{2}_{2}$
$\hskip 6.25963pt+\|v_{i|k}-v_{\textup{ref}}\|^{2}_{2}$
s.t. $\forall i=0,\dots,29$
$x_{i+1|k}=f(x_{i|k},u{i|k})$
$u_{i|k}\in\mathcal{U}$
$x_{0|k}=x(k)$
This implementation aims to reach the opposite side of the initial condition of the area, while keeping the velocity constant at $v_{\textup{ref}}=3\textup{ms}^{-1}$. Notably, the optimisation problem is not aware of the obstacles, such that obstacle avoidance is purely a result of the safety filter. The closed-loop results for $5$ initial conditions over $70$ time steps are shown in Figure 3, where the system successfully avoids the obstacles in the cluttered environment, even at the high initial speed of $3\textup{ms}^{-1}$. The solve time for the neural network inference for safety verification, i.e. lines 4 and 5 in Algorithm 1, take on average $0.26$ms and solving the two optimisation problems takes on average $37$ms. Comparatively, the PCBF algorithm in [12] requires on average $500$ms to solve the two predictive optimisation problems. The computational speedup is therefore a factor of $13$ in the case of overwriting the inputs and $1900$ when the proposed input is deemed safe. Finally, we provide the magnitude of the safety interventions for one of the trajectories in Figure 4. It can be seen that the safety filter not only overrides the steer change before an obstacle is hit, but also the acceleration such that the car is able to navigate around the obstacles safely.
Figure 3: Closed-loop trajectories of the simulated car dynamics, where multiple initial conditions with a velocity of $3$m/s and an initial steering angle of $\delta=0$ are shown (car). The performance controller is given by a naive MPC implementation which is not aware of the obstacles, such that the distance to the opposite side of the area is minimised. States where the safety filter intervened are marked in red.
Figure 4: Magnitude of safety filter interventions of the steer change (top) and the acceleration (bottom) for the trajectory with initial position $(-1.8,-1.0)$.
7 CONCLUSION
We propose a computationally efficient safety filter formulation which relies on approximating the optimal value function of a predictive control method. This function encodes the information regarding safety in terms of state constraint satisfaction with respect to the entire prediction horizon of the predictive optimisation problem and enhances the safe set and domain of attraction of a terminal control barrier function. By using the explicit approximation, a horizon independent optimisation problem can be formulated with a single decrease constraint of the control barrier function. We advance the theoretical properties with respect to exogeneous disturbances of the original algorithm in [12] and translate the results to input-to-state stability of the proposed algorithm with respect to approximation errors. Finally, we propose a negative continuous extension of the predictive control barrier function in order to mitigate approximation errors within the safe set. The theoretical properties are validated in an illustrative linear example and the practicality of a computationally advantageous safety filter is shown in a miniature race car simulation.
In order to achieve a computationally efficient sampling of the extended PCBF $h$ in (21), we propose a modification of the geometric sampling method presented in [41] in Algorithm 2. The geometric sampling method allows using previously computed solutions to the predictive optimisation problems for the computation of $h$ as a warmstart as well as storing information if the previous sample on a line search was within $\mathcal{S}_{\mathrm{PB}}$ or $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, reducing the number of times two optimisation problems have to be solved.
Algorithm 2 Sampling $h$
1:  Generate goal points $\mathbb{G}$ according to [41, Algorithm 2]
2:  Initialise boolean $\mathrm{in\_}\mathcal{S}_{\mathrm{PB}}\leftarrow\mathrm{False}$
3:  Choose initial seed $\mathbb{S}\leftarrow\{\mathbf{s}\}=\{(x,\mathbf{o},\mathrm{in\_}\mathcal{S}_{\mathrm{PB}})\}$
4:  Initialise data set $\mathbb{D}\leftarrow\emptyset$
5:  Choose step size $d>0$
6:  for goal point $g\in\mathbb{G}$ (in parallel) do
7:     Sample seed $\mathbf{s}_{0}\leftarrow(x_{0},\mathbf{o}_{0},\mathrm{in\_}\mathcal{S}_{\mathrm{PB},0})\in\mathbb{S}$
8:     Compute number of points on line $n\leftarrow\|x_{0}-g\|/d$
9:     for $i=1,\dots,n$ do
10:        $x_{i}\leftarrow x_{0}+i(g-x_{0})/n$
11:        if $\mathrm{in\_}\mathcal{S}_{\mathrm{PB},i-1}$ then
12:           Obtain $\mathbf{o}_{i}$, $h(x_{i})$ from (22) with warmstart $\mathbf{o}_{i-1}$
13:           if infeasible then
14:              $\mathrm{in\_}\mathcal{S}_{\mathrm{PB},i}\leftarrow$ False
15:              Obtain $\mathbf{o}_{i}$, $h(x_{i})$ from (10) with warmstart $\mathbf{o}_{i-1}$
16:           else
17:              $\mathrm{in\_}\mathcal{S}_{\mathrm{PB},i}\leftarrow$ True
18:           end if
19:           $\mathbf{s}_{i}\leftarrow(x_{i},\mathbf{o}_{i},\mathrm{in\_}\mathcal{S}_{\mathrm{PB},i})$
20:           $\mathbb{D}\leftarrow\mathbb{D}\cup(x_{i},h(x_{i}))$
21:        else
22:           Obtain $\mathbf{o}_{i}$, $h(x_{i})$ from (10) with warmstart $\mathbf{o}_{i-1}$
23:           if $h_{\mathrm{PB}}(x_{i})=0$ then
24:              $\mathrm{in\_}\mathcal{S}_{\mathrm{PB},i}\leftarrow$ True
25:              Obtain $\mathbf{o}_{i}$, $h(x_{i})$ from (22) with warmstart $\mathbf{o}_{i-1}$
26:           else
27:              $\mathrm{in\_}\mathcal{S}_{\mathrm{PB},i}\leftarrow$ False
28:           end if
29:           $\mathbf{s}_{i}\leftarrow\{(x_{i},\mathbf{o}_{i},\mathrm{in\_}\mathcal{S}_{\mathrm{PB},i})\}$
30:           $\mathbb{D}\leftarrow\mathbb{D}\cup(x_{i},h(x_{i}))$
31:        end if
32:     end for
33:     $\mathbb{S}\leftarrow\mathbb{S}\cup\mathbf{s}_{n}$
34:  end for